A fish tank with a rectangular base has a volume of 3,360 cubic inches. The length and width of the tank are 14 inches and 12 inches, respectively. Find the height, in inches, of the tank.
1
Expert's answer
2012-06-05T08:00:01-0400
For a tank with rectangular base the volume is calculated as V=L*W*H where H-height, L-length, W-width.
Therefore H=V/L/W
Substituting given values we obtain
H=3 360/14/12=20 inches